OpenStudy (anonymous):

CONVERT 1/8 TO A DECIMAL AND A PERCENTAGE

OpenStudy (anonymous):

1/8=0.125 and 0.125/100= 0.00125%

OpenStudy (anonymous):

Actually, you would need to multiply the decimal by hundred to get the percentage. So 1/8=0.125 =12.5%.